Lynne jogs daily around a circular track of diameter 25 m for a total distance of


942 m in 5.0 min. (a) How many times did Lynne jog around the track? (b) What


is Lynne’s average angular speed in radians/s?

1
Expert's answer
2021-11-07T19:26:29-0500

1) Let us calculate the length of the circular path with diameter 25 m

"L=\pi D \approx 78.54\,\mathrm{m}."

The number of turns is

"N= \dfrac{S} {L} = \dfrac{942}{78.54} \approx 12."


2) The average angular speed in radians per second is

"\omega =\dfrac{2\pi}{T}," where T is time needed for jogging around one circular track.

"T =\dfrac{5\cdot 60\,\mathrm{s}}{N} = 25\,\mathrm{s}."

Therefore, the angular velocity is

"\omega=\dfrac{2\pi}{25\,\mathrm{s}} \approx 0.25\,\mathrm{rad/s}."